INAF queue observing, 2018 November 11-12
Observer: R. Carini, M. Faccini
Telescope Operator: D. Gonzal-Huerta
Support Astronomer: O. Kuhn
Instrument: LUCI1 (LS mode), LUCI2
Night duration (12\xB0 el twilight) 1:15-12:45 UT = 11h 30m
Observing time: 10h 45min (93%)
Weather time loss: h 45m (7%)
technical time loss 0h m (%)
At the sunset the sky is mostly clear, wind with gusts around 20 m/s, humidity about 20.5%
|| ExpTime /Time_tot
|| Cresci a1, Saracco
|| Luci1==2, luci2
|| Mods1 + MOds2
|| dual grating, blue grating
|| completed, some images are not recovered during the transit.
| Nisini telluric
|| luci2 mos
|| G210 HiRes, K,J,H
|| field 4
|| luci2 mos
|| G210 HiRes, K,J,H
- flat+arc Cresci A1
- Saracco, flatarc luci2
TO opens the dome. the deley is due to burst of wind.
collimation on the target.
TO closes the dome, gusts of wind up to 22 m/s
we open again
- 02:20 problem with the collimation
start the acquisition of the target of Mancini Hat-P-47
- seeing about 1.5"
- seeing variable.
- 03:38 seeing up to 2.0"
- fitsflush red (mods1), fistflus blue (mods2). Some images are not recovered.
- seeing more than 2.0"
- 04:46 seeing 1.2"-1.8"
- 04:53 seeinf0.9"-1.7"
- 4:59* restart the script with mods1-red
- seeing 1.5'-20''
- 5:28* restart the script with mods2-blue
- seeing 1,5" - 2,0"
- 06:06* restart the sript with mods1-blue
- seeing 1.1" - 1.8"
- 06:21 elevation 87:19 over the limits for the guiding of the telescope. Now we are in tracking mode and it is possible to loose the target.
- 06:28 we transit the meridian it seems we haven' lost data till now
- 06:35 restart guiding. Apparently no data lost
- 07:08 MODS1 BLue CCCD temperature of -11 C exceede -90C -detector warming up!
- 07:09 MODS1 Blue CCD temperature is back normale range
- 07:10 seeing 1 - 1,2"
- 07:10 restart MOds1 red grating
- 07:35 stop the integration
slew to Moretti program, target A6
- we abort and take an image to verify the the star is already in the slit
- seeing about 1.0"
- 08:00 seeing jump to 1.7" and decreases to 1.1"
- seeing very variable up to 1.8", but the mean seeing is 1.2"
problem with the observin procedure.
* We start again the acquisition, offset failed
- we start again the acquisition error during the preset, we send again the script
- 09:17 error again
- 09:18 try again the preset and it failed
- The problem dipend on the script AcQMOds that does not ignore "syncoffset" when only MODS1 or MODS2 was used, even if it should do it.
slew to De Carli program. we try the new procedure, but we are not sure that it is going to work.
- we try to acquire the field in the mos using just a slit, as in the case of the telluric with LUCI, but it failes.
- ERROR: Slit mask ID592104 is a multi-object mask!
- You must perform a MOS mask alignemtn using an mms. ModsAlign aborting
switch tu LUCI
telluric nisini field 4
slew to scientific target, field4
luci2.0048-0069 0069 not good
- seeing about 1.3"
- 11:06 seeing 1.4"
- 11:21 seeing 1.5"
- 11:32 seeing 1.4"
- 11:34 seeing 1.8"
- 11:42 seeing 1.4"
- 12:02 seeing 1.2"
stop the integration, airmass 2.00, the images are good but not the last
* slew to Antoniucci target V900 mon
* 12;25 check pointing
send again the script
- sci HKSpec
telluric in HKspec
- acq ==luci2.0078-0080
- bias Mancini
- cal wave mods1 and mods2
- mos flat
- send again mos flat mods2
flat arc Nisini field 4 H